OpenStudy (anonymous):

Simplify the expression: tan(sin^−1(x))

OpenStudy (john_es):

This expression means, "calculate the tangent of the angle whose sin is x". So, we know by trigonometric identities, \[\sin^2\alpha+\cos^2\alpha=1\]\[1+\tan^2\alpha=\frac{1}{\cos^2\alpha} \Rightarrow\tan^2\alpha=\frac{1}{1-\sin^2\alpha} -1\] with \[\alpha=\sin^{-1}x\]Or \[\sin\alpha=x\]So \[\tan^2\alpha=\frac{1}{1-x^2} -1=\frac{x^2}{1-x^2}\]\[\tan(\sin^{-1}x)=\pm\frac{x}{\sqrt{1-x^2}}\]